OpenStudy (anonymous):

remove parentheses and simplify 8c+5d-6(3c-2d)=?

OpenStudy (anonymous):

you can remove the parentheses by simply distributing the -6 into the parenthese then collect like terms

OpenStudy (anonymous):

I am still clueless lol

OpenStudy (anonymous):

8c+5d-6(3c-2d)=? 8c+5d-(18c-12d)=? 8c+5d-18c+12d=17c-10d the answer is 17c-10d...=)

OpenStudy (gw2011):

-6(3c-2d)=-18c+12d Now you get the following: 8c+5d-18c+12d=-10c+17d

OpenStudy (anonymous):

oops.....the variables are interchanged in the final answer...sorry...=)